### v3

@@ 3.3 Model Architecture
To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [32]. Overall, the position loss is defined as
@@
By employing the Hungarian matching algorithm [16] to optimize Eq. 1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates. Notice that the matching cost between a repetitive action instance and $\emptyset$ doesn’t depend on the prediction, which means that in that case the cost is a constant. Following this step, we can apply the corresponding losses related to classification and temporal location prediction.
We define a Hungarian loss for all pairs matched in the previous step to supervise both classification and temporal location predictions, i.e. a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq. 3:
@@
where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq. 1.
@@ 3.4 Dynamic Action Query
As discussed in Sec. 1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined. Therefore, we propose the Dynamic Action Query strategy, which adaptively updates the action query $\mathbf{Q}^{\text{act}}$ using the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}=\mathbf{E}^{\text{act}}$, as shown in Fig. 2. The DAQ strategy, simple yet effective, provides priors for the decoder to dynamically and contextually attend to “repetitive actions” based on the input video content, eliminating the need for manually defining action categories and thereby enhancing the model’s generalization capability. This mechanism ensures that the cross-attention module in the decoder pools features from the entire video while focusing on relevant local repetitive features. Such dynamic adjustment allows the model to concentrate on pertinent “repetitive actions”, offering a deeper understanding of action queries and improving training convergence.
In summary, the DAQ approach not only circumvents the need for manual action category definitions but also leverages dynamic, contextually aware priors to enhance both the focus and generalization capabilities of the model. Our supplementary material further explores various methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$, confirming that the DAQ strategy is the most effective.
@@ 3.5 Inter-query Contrastive Learning
Since the input video may contain other distractors such as the background motion  [^2] We assume that each video contains only a primary repetitive action type, which is also the case for the public dataset., another unique challenge to the TRC task is to recognize action instances with identical content. This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations. To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries. Intuitively, based on the classification predictions $\mathbf{A}$, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ inter-query contrastive learning on them, as shown in Fig. 2 (right). The features that are classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$. Then we apply contrastive learning using InfoNCE loss [12] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
@@ 3.6 Training
where $\lambda_{\text{Hungarian}},\lambda_{\text{ctrs}}\in\mathbb{R}$ are the coefficients. Following DETR [3], we also found it helpful to use auxiliary losses in the decoder during training. Specifically, we add $\mathcal{L}_{\text{Hungarian}}$ loss on the predictions from the prediction heads right after the encoder i.e. $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{P}}$, and add prediction heads and $\mathcal{L}_{\text{Hungarian}}$ loss after each decoder layer. All prediction heads share their parameters. This ensures the model focuses on the correct features at each stage consistently, thereby accelerating convergence.
@@ 4.1 Datasets and Metrics
RepCountA dataset [14] is currently the largest and most challenging benchmark for the video TRC task  [^3] The RepCountB [14] test subset is proprietary and not publicly available.. It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments. This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [9, 33, 18, 47]. We use the start and end positions of each action instance to compute the GT label $\hat{\mathbf{P}}$ provided by the annotations. We train our model on the RepCountA train set and select the best model on the validation set. We report the evaluation results on the test set.
Table 1: Comparison to the state-of-the-arts on RepCountA [14] dataset. We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with $\dagger$). We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
@@
X3D [10] X3D 0.9105 0.1059 - - - - - -
TANet [26] TANet 0.6624 0.0993 - - - - - -
VideoSwinT [25] ViT 0.5756 0.1324 - - - - - -
GTRM [15] I3D 0.5267 0.1589 - - - - - -
RepNet [9] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
Zhang et al. [47] 3D-ResNext 0.8786 0.1554 - - - - - -
TransRAC [14] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
Li et al. [19] ViT 0.3841 0.3860 - - - - - -
TadTR [24] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
ActionFormer [45] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
ReAct [34] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
@@
UCFRep dataset [47] is a subset of the UCF101 dataset [36], including fitness videos and daily life videos. Following previous work [14, 19], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
Metrics. Following previous works [14, 9, 47, 19], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance. OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count ±$\pm 1$ range. MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts. Formally,
@@ 4.2 Implementation Details
We implement our approach with different backbone video feature extractors respectively, including TSN [41], I3D [4], and ViT [8]. For encoder-decoder Transformer, we set ${L}^{\text{enc}}=2$ and ${L}^{\text{dec}}=4$, both with 8-head attention mechanisms. The feature dimension is set to $C=512$. We set $Q=40$ queries in our method empirically. For more results with different feature channels $C$ and query numbers $Q$, please refer to the supplementary materials. The length of the video input is set to $T=512$ frames without down-sampling. We utilize the AdamW optimizer [27] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs. We set $\lambda_{\text{Hungarian}}=1.0$, $\lambda_{\text{L1}}=5.0$, $\lambda_{\text{gIoU}}=0.4$, $\lambda_{\text{ctrs}}=1.0$, confidence threshold $\alpha=0.2$. We provide further implementation details in the supplementary.
@@ 4.3 Comparison to State-of-the-arts
Table 2: Generalization comparison with SOTA TRC methods on UCFRep [47] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
@@
RepNet [9] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
TransRAC [14] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
Li et al. [19] 3D-ResNext 0.5227 0.3500 - - - - - -
@@
Table 3: Comparison of computational complexity. We benchmark all the methods with inference on the RepCountA dataset.
       $T$        Params (M)        FLOPs (G)
       RepNet [9]        64        20.47        92.40
       TransRAC [14]        64        14.43        100.19
       Ours        64        18.87        1.44
       Ours        512        18.87        6.06
Results on RepCountA dataset. We compare our proposed approach to the state-of-the-art methods on the RepCountA [14] dataset following previous work [14, 19] in Tab. 1. We compare with the SOTA action recognition [10, 26, 25], action segmentation [15] methods (top block), and TRC [9, 47, 14, 19] approaches (second block). We further adapt recent DETR-style action detection approaches [24, 45, 34] for the TRC task (third block). We change their output layers accordingly and train them on RepCountA.
As shown in Tab. 1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [9, 14, 19] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [24, 45, 34] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not support them to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
Results on UCFRep dataset. We also evaluate the generalization ability of our method. Following previous work [14, 19], we evaluate the model trained on the RepCountA dataset [14] on UCFRep [47] test set. Tab. 2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions. The results demonstrate the effectiveness and generalization capability of our method.
@@
Figure 4: Visualization of failure case on RepCountA [14] dataset. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts. Each colored block represents a GT or predicted action instance.
Figure 5: Qualitative results on UCFRep dataset. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map. The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [47].
Table 4: Effect of DAQ and ICL modules on RepCountA [14] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
@@
Efficiency. Additionally, we evaluate the computational complexity of our method in comparison to state-of-the-art methods. We benchmark all the methods with inference on the RepCountA dataset. For fairness of comparison, we calculate the Floating Point Operations (FLOPs) in the counting module for all the methods, excluding the video feature extraction process. We set the batch size to be $1$ and clip length to be $64$ for each method. Tab. 3 shows that our query-based method significantly reduces the computation load compared to the previous similarity-matric-based methods [9, 14], while also delivering superior accuracy. Importantly, the difference in computational demand is expected to widen with longer input clip lengths. This is because the computational complexity for similarity-matrix-based methods increases quadratically, whereas it grows linearly in our approach. Unfortunately, due to computational limitations, we are unable to obtain the results of RepNet [9] and TransRAC [14] when $T=512$.
@@ 4.4 Qualitative Results
We visualize the predictions of our approach and baseline method TransRAC [14] on the RepCountA [14] dataset in Fig. 3. TransRAC [14] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map. However, this approach suffers from a lack of interpretability and finally results in miscounting. In contrast, our approach not only produces an accurate final count but also correctly localizes the action start and end positions (colored blocks) in most cases. In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig. 3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline. Please refer to our supplementary video for more qualitative results. We present the video result of the first case in Fig. 3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving. Our method selectively counts only the pull-up actions, as they are the primary focus and consistent with the main action instances of the video, excluding dissimilar actions like “waving”, validating the robustness and generalization ability of our method.
In addition, our query-based representation offers excellent interpretability, making it easy to identify the issues in case of failures. For example, Fig. 4 shows a typical failure case. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
We further illustrate the generalization performance of the proposed method on the unseen UCFRep [47] test set in Fig. 5. We directly apply our trained model and do not use UCFRep training data. Our model still accurately recognizes the action instances and gets the correct count in the challenging cases, indicating robust generalization ability. Specifically, the top case exhibits extreme viewpoint and lighting conditions, while the bottom case contains the action of soccer juggling which is not seen in the training set. We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions.

### v4

@@ 3.3 Model Architecture
To measure the similarity of the predicted temporal locations and the GT temporal location, we define $\mathcal{L}_{\text{pos}}$ using the linear combination of the $L_{1}$ distance and the generalized Intersection over Union (IoU) loss [rezatofighi2019generalized]. Overall, the position loss is defined as
@@
By employing the Hungarian matching algorithm [kuhn1955hungarian] to optimize Eq.˜1, we can achieve the final optimal matching $\hat{\sigma}$ which uniquely assigns a prediction to a GT action cycle, finding one-to-one matching without duplicates. Notice that the matching cost between a repetitive action instance and $\emptyset$ doesn’t depend on the prediction, which means that in that case the cost is a constant. Following this step, we can apply the corresponding losses related to classification and temporal location prediction.
We define a Hungarian loss for all pairs matched in the previous step to supervise both classification and temporal location predictions, i.e. a linear combination of a negative log-likelihood loss for class prediction and a temporal position loss defined in Eq.˜3:
@@
where $\hat{\sigma}$ is the optimal assignment computed in the previous bipartite matching step, i.e. Eq.˜1.
@@ 3.4 Dynamic Action Query
As discussed in Sec.˜1, the TRC problem requires recognizing open-set action instances depending on the video content, where the action category is not predefined. Therefore, we formulate the problem as a binary classification task with two categories: “repetitive actions” and “others ($\emptyset$)”. The key idea is that the class of “repetitive actions” is not fixed or manually defined. Rather, it is dynamically inferred based on the video content. To achieve this, we propose the Dynamic Action Query strategy. As shown in Fig.˜2 (top right), DAQ adaptively updates the action query $\mathbf{Q}^{\text{act}}$ by directly assigning it the selected encoder action query features $\mathbf{E}^{\text{act}}$, i.e. $\mathbf{Q}^{\text{act}}\leftarrow\mathbf{E}^{\text{act}}$. These decoder action queries $\mathbf{Q}^{\text{act}}$ embed the video content, enabling the decoder to dynamically define “repetitive actions” based on the input video content. The DAQ strategy, simple yet effective, eliminates the need for manually defining action categories and thereby enhances the model’s generalization capability.
Besides, we also explore several different methods for initializing the queries in the decoder $\mathcal{D}(\cdot)$ in the supplementary, confirming that the DAQ strategy is the most effective.
@@ 3.5 Inter-query Contrastive Learning
Since the input video may contain other distractors such as the background motion [^2] We assume that each video contains only a primary repetitive action type, which is also the case for the public dataset., another unique challenge to the TRC task is to recognize action instances with identical content. This requires that the actions of interest we classify exhibit similarity in their motion patterns, while other action queries should have dissimilar representations. To tackle this challenge, we propose Inter-query Contrastive Learning to distinguish the action queries, as shown in Fig.˜2 (right). ICL encourages the model to pull together queries corresponding to the same repetitive actions in the feature space, while pushing apart those related to background or unrelated movements. Therefore, we partition the action queries decoded by the decoder $\mathbf{D}^{\text{act}}$ into two categories and employ contrastive learning on them based on the classification predictions $\mathbf{A}$. Specifically, features classified as “repetitive actions” form the positive set $S^{+}$, while the other features form the negative set $S^{-}$. Then we apply contrastive learning using InfoNCE loss [he2020momentum] $\mathcal{L}_{\text{ctrs}}$ over the representation space:
@@ 3.6 Training
where $\lambda_{\text{Hungarian}},\lambda_{\text{ctrs}}\in\mathbb{R}$ are the coefficients. Following DETR [carion2020end], we also found it helpful to use auxiliary losses in the decoder during training. Specifically, we add $\mathcal{L}_{\text{Hungarian}}$ loss on the predictions from the prediction heads right after the encoder i.e. $\tilde{\mathbf{A}}$ and $\tilde{\mathbf{P}}$, and add prediction heads and $\mathcal{L}_{\text{Hungarian}}$ loss after each decoder layer. All prediction heads share their parameters. This ensures the model focuses on the correct features at each stage consistently, thereby accelerating convergence.
@@ 4.1 Datasets and Metrics
RepCountA dataset [hu2022transrac] is currently the largest and most challenging benchmark for the video TRC task  [^3] The RepCountB [hu2022transrac] test subset is proprietary and not publicly available.. It is primarily compiled from fitness videos on YouTube, including a wide range of fitness activities conducted in diverse settings, including homes, gyms, and outdoor environments. This dataset stands out due to its extensive video lengths, significant variations in the average motion cycle, and a higher number of repetitive cycles compared to prior datasets [dwibedi2020counting, runia2018real, levy2015live, zhang2020context]. We use the start and end positions of each action instance to compute the GT label $\hat{\mathbf{P}}$ provided by the annotations. We train our model on the RepCountA train set and select the best model on the validation set. We report the evaluation results on the test set.
Table 1: Comparison to the state-of-the-arts on RepCountA [hu2022transrac] dataset. We compare with SOTA action recognition/segmentation methods (top block), TRC methods (second block), and action detection methods (third block with $\dagger$). We further report MAE and OBO metrics for short-, medium-, and long-period test actions.
@@
 X3D [feichtenhofer2020x3d] X3D 0.9105 0.1059 - - - - - -
TANet [liu2021tam] TANet 0.6624 0.0993 - - - - - -
VideoSwinT [liu2022video] ViT 0.5756 0.1324 - - - - - -
GTRM [huang2020improving] I3D 0.5267 0.1589 - - - - - -
 RepNet [dwibedi2020counting] ResNet 0.5865 0.2450 0.7793 0.0930 0.5893 0.1591 0.4549 0.4062
Zhang et al. [zhang2020context] 3D-ResNext 0.8786 0.1554 - - - - - -
TransRAC [hu2022transrac] ViT 0.4891 0.2781 0.5789 0.0233 0.4696 0.2955 0.4420 0.4375
Li et al. [li2024repetitive] ViT 0.3841 0.3860 - - - - - -
 TadTR [liu2022end] $\dagger$ I3D 1.1314 0.0662 0.8364 0.0233 1.1591 0.0000 1.3106 0.1406
ActionFormer [zhang2022actionformer] $\dagger$ I3D 0.4990 0.2781 0.4164 0.1628 0.3768 0.3409 0.6385 0.3125
ReAct [shi2022react] $\dagger$ TSN 0.4592 0.3509 0.2805 0.1576 0.3037 0.4318 0.6862 0.3906
@@
UCFRep dataset [zhang2020context] is a subset of the UCF101 dataset [soomro2012ucf101], including fitness videos and daily life videos. Following previous work [hu2022transrac, li2024repetitive], we do not use the train set but directly test our model on the test set to evaluate the model generalization ability.
Metrics. Following previous works [hu2022transrac, dwibedi2020counting, zhang2020context, li2024repetitive], we compute two commonly used metrics, OBO and MAE, to evaluate the model performance. OBO (Off-By-One count error) measures the percentage that the predicted count is within the GT count $\pm 1$ range. MAE (Mean Absolute Error) measures the normalized absolute difference between the predicted and GT counts. Formally,
@@ 4.2 Implementation Details
We implement our approach with different backbone video feature extractors respectively, including TSN [wang2016temporal], I3D [carreira2017quo], and ViT [dosovitskiy2021an]. For encoder-decoder Transformer, we set ${L}^{\text{enc}}=2$ and ${L}^{\text{dec}}=4$, both with 8-head attention mechanisms. The feature dimension is set to $C=512$. We set $Q=40$ queries in our method empirically. For more results with different feature channels $C$ and query numbers $Q$, please refer to the supplementary materials. The length of the video input is set to $T=512$ frames without down-sampling. We utilize the AdamW optimizer [loshchilov2017decoupled] with a learning rate of $0.002$, a batch size of $64$, and train the model for $80$ epochs. We set $\lambda_{\text{Hungarian}}=1.0$, $\lambda_{\text{L1}}=5.0$, $\lambda_{\text{gIoU}}=0.4$, $\lambda_{\text{ctrs}}=1.0$, confidence threshold $\alpha=0.2$. We provide further implementation details in the supplementary.
@@ 4.3 Comparison to State-of-the-arts
Table 2: Generalization comparison with SOTA TRC methods on UCFRep [zhang2020context] dataset. MAE and OBO metrics for short-, medium-, and long-period actions are also reported.
@@
 RepNet [dwibedi2020counting] ResNet 0.5336 0.2984 0.6219 0.1739 0.4825 0.3600 0.4996 0.5000
TransRAC [hu2022transrac] ViT 0.6180 0.3143 0.6296 0.1951 0.5842 0.4250 0.6784 0.4118
Li et al. [li2024repetitive] 3D-ResNext 0.5227 0.3500 - - - - - -
@@
Table 3: Comparison of computational complexity and inference time across different models. We evaluate the complexity and efficiency on varying frame lengths ($T$). OOM denotes “Out-of-Memory”, indicating that the model ran out of memory on a 32G GPU, and therefore, no inference time could be reported.
    Model    $T$    Params (M)    FLOPs (G)    Inference Time (s)
    RepNet [dwibedi2020counting]    64    43.98    163.44    3.62
   512    48.94    6832.50    OOM
   TransRAC [hu2022transrac]    64    42.28    582.02    2.43
   512    48.22    8555.78    OOM
        64    42.38    72.48    0.15
   Ours    512    42.38    574.38    0.78
Results on RepCountA dataset. We compare our proposed approach to the state-of-the-art methods on the RepCountA [hu2022transrac] dataset following previous work [hu2022transrac, li2024repetitive] in Tab.˜1. We compare with the SOTA action recognition [feichtenhofer2020x3d, liu2021tam, liu2022video], action segmentation [huang2020improving] methods (top block), and TRC [dwibedi2020counting, zhang2020context, hu2022transrac, li2024repetitive] approaches (second block). We further adapt recent DETR-style action detection approaches [liu2022end, zhang2022actionformer, shi2022react] for the TRC task (third block). We change their output layers accordingly and train them on RepCountA.
As shown in Tab.˜1, our approach significantly outperforms the state-of-the-art methods across actions of varying lengths. Specifically, the similarity-matrix-based methods [dwibedi2020counting, hu2022transrac, li2024repetitive] suffer from quadratic computation complexity. To manage this, they employ a sparse sampling strategy, ensuring reasonable content coverage within a limited temporal context window. However, this approach results in inferior performance for short, rapid action instances, as it tends to overlook the cycles. Conversely, the action detection approaches [liu2022end, zhang2022actionformer, shi2022react] are primarily developed for detecting action instances specific to particular classes. To adapt them to the TRC task, we modify their output layer to incorporate class-agnostic supervision accordingly. However, relying solely on class-agnostic supervision does not support them to dynamically identify repetitive cycles based on the input videos. As a result, their performance on long, slow action instances is generally inferior compared to dedicated TRC approaches. In contrast, our approach effectively balances the detection of actions at various speeds.
Results on UCFRep dataset. We also evaluate the generalization ability of our method. Following previous work [hu2022transrac, li2024repetitive], we evaluate the model trained on the RepCountA dataset [hu2022transrac] on UCFRep [zhang2020context] test set. Tab.˜2 shows that our approach generally outperforms existing works, with more significant improvements observed in longer-period actions. The results demonstrate the effectiveness and generalization capability of our method.
@@
Figure 4: Visualization of failure case on RepCountA [hu2022transrac] dataset. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts. Each colored block represents a GT or predicted action instance.
Figure 5: Qualitative results on UCFRep dataset. Each colored block represents a GT or predicted action instance. TransRAC represents the results by density map, and the final count value is obtained by summing the values in the density map. The vertical lines in the labels represent the time points at which the actions begin since only the starting point annotations are provided in UCFRep [zhang2020context].
Table 4: Effect of DAQ and ICL modules on RepCountA [hu2022transrac] dataset. (a) ablates DAQ strategy. (b) ablates the ICL strategy. (c) is our full model.
@@
Efficiency. Additionally, we evaluate the computational complexity of our method in comparison to state-of-the-art methods. We benchmark all the methods with inference on the RepCountA dataset. As shown in Tab.˜3, we present a detailed comparison including parameter count, FLOPs, and inference time, all measured on a single NVIDIA V100 32G GPU with a bacth size to be 1. Each method’s performance across two varying frame lengths ($T=64$ and $T=512$) is evaluated on the full model, to provide a more comprehensive assessment.
Notably, prior methods like RepNet [dwibedi2020counting] and TransRAC [hu2022transrac] were originally designed to accept video inputs primarily at $T=64$ frames. When we attempted to process $T=512$ frames with their implementations, their computational complexity increased drastically. This is because both methods rely on a temporal similarity correlation matrix, causing their complexity to grow quadratically with the input length $T$. Consequently, they encountered “Out-of-Memory (OOM)” errors during execution, failing to obtain their inference time.
In contrast, our method is significantly more lightweight and achieves demonstrably faster inference times compared to prior approaches. More importantly, its computational complexity scales linearly with increasing input video length ($T$), which robustly ensures the algorithm’s efficiency and allows it to maintain a very fast processing speed even when processing $T=512$ frames, as evidenced in Tab.˜3.
Please refer to Supplementary for a decomposed model component complexity analysis.
@@ 4.4 Qualitative Results
We visualize the predictions of our approach and baseline method TransRAC [hu2022transrac] on the RepCountA [hu2022transrac] dataset in Fig.˜3. TransRAC [hu2022transrac] represents action cycles using density maps, and the final count value is obtained by summing the values in the density map. However, this approach suffers from a lack of interpretability and finally results in miscounting. In contrast, our approach not only produces an accurate final count but also correctly localizes the action start and end positions (colored blocks) in most cases. In addition, our method exhibits robustness to changes in viewpoint, background noise, and sudden interruptions, e.g., in the second case of Fig.˜3, we accurately estimate the time positions even with viewpoint changes as well as a sudden interruption in the middle of the timeline. Please refer to our supplementary video for more qualitative results. We present the video result of the first case in Fig.˜3, which shows that while the subject performs pull-ups, there are moments of talking and arm waving. Our method selectively counts only the pull-up actions, as they are the primary focus and consistent with the main action instances of the video, excluding dissimilar actions like “waving”, validating the robustness and generalization ability of our method.
In addition, our query-based representation offers excellent interpretability, making it easy to identify the issues in case of failures. For example, Fig.˜4 shows a typical failure case. Due to the excessive zooming in, the legs of the human body are truncated, making a large difference in the action motion feature, and resulting in several missed cycle counts.
We further illustrate the generalization performance of the proposed method on the unseen UCFRep [zhang2020context] test set in Fig.˜5. We directly apply our trained model and do not use UCFRep training data. Our model still accurately recognizes the action instances and gets the correct count in the challenging cases, indicating robust generalization ability. Specifically, the top case exhibits extreme viewpoint and lighting conditions, while the bottom case contains the action of soccer juggling which is not seen in the training set. We attribute the performance advantage to the proposed DAQ and ICL designs, which empower the model to adaptively adjust the action queries based on the input video features and effectively localize similar (repetitive) action instances, distinguishing them from the background noise actions.